What is the equation of a circle touching the lines x-3y-11=0 and 3x-y-9=0 and having its center on the line x-2y+19=0 ?

Answer by Alan3354(69443)

What is the equation of a circle touching the lines x-3y-11=0 and 3x-y-9=0 and having its center on the line x-2y+19=0 ?
----------------------
Find the bisector of the 2 lines that are tangent.
The intersection of the bisector and x-2y+19=0 is the center of the circle.
The radius of the circle is the distance from the center to either of the tangent lines.
--------------
This is a multi-step problem. I've done it once or twice.
email via the TY note for help or to check your work.